# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

### v1

@@ 5.1 Scalability
Figure 7 and Figure 7 hows the calculation times across the various network sizes of the synthetic data and IBM datasets, respectively. The time in Figure 7 is shown on a logarithmic scale since this provides a better distinction among the methods across dataset sizes. These results indicate that FlowScope is the fastest of all methods, although GARG-AML seems to scale better to larger datasets, as the calculation time increases much more slowly. In contrast, AutoAudit—although quite fast in the beginning—has difficulties with networks of the largest size. In addition, the distribution of the calculation time of AutoAudit is much wider and has more outliers, which indicates that it is highly sensitive to the network topology.
@@ 5.2.2 IBM Data
This section presents the results on the IBM datasets for the smurfing patterns, i.e., gather-scatter and scatter-gather, and the full ‘Is Laundering’ label. The thresholds applied for binary label definition are 0.1, 0.2 and 0.9. The results for the HI-Small and LI-Large datasets are presented in Figure 12 and Figure 13, respectively. The results for all patterns and thresholds are provided in Appendix A.
@@ 6.1 Limitations in the Data
GARG-AML assumes full visibility into all transactions, which is rarely the case in practice [3, 40, 13]. Banks typically have access only to transactions that involve their own clients, including sender and receiver account numbers and client attributes. Transactions between two external accounts remain hidden, which makes second-order neighbourhoods incomplete. This can obscure smurfing structures or lead to incorrect classification (see Figure 14). Moreover, banks cannot link multiple third-party accounts that are owned by the same individual, thus further weakening pattern recognition.
Figure 14: Illustration of smurfing over different banks, which are represented by different colours. The (i)-icon indicates that the main (blue) bank has all client information about the corresponding account. The dotted lines indicate that the bank is unaware of these transactions/accounts.
@@ 7 Conclusion
This paper introduces Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), which is a novel approach for detecting smurfing by quantifying the structural similarity of nodes’ neighbourhoods to known money laundering patterns via density-based features in the adjacency matrix.
We validated GARG-AML on several synthetic datasets and compared it to state-of-the-art smurfing detection methods. The results of the experiments show that GARG-AML achieves strong detection power, with state-of-the-art performance of GARG-AML scores, and excellent scalability for large networks. Unlike approaches that are based on complex optimization techniques or machine learning models, GARG-AML relies on domain expertise and straightforward density-based calculations.
@@
Detecting smurfs is also an important research area for network security and intrusion detection [49, 38]. Further research could analyse whether similar featurisation of the patterns is possible for denial of service (DoS) or similar attacks.
Acknowledgments
This work was supported by the Research Foundation – Flanders (FWO) [grant number 1SHEN24N] and by the BNP Paribas Fortis Chair in Fraud Analytics. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
Author Contributions
Bruno Deprez: Conceptualization, Data Curation, Methodology, Formal Analysis, Software, Visualization, Writing—Original Draft.
Bart Baesens: Conceptualization, Supervision, Writing—Review & Editing.
Tim Verdonck: Conceptualization, Supervision, Writing—Review & Editing.
Wouter Verbeke: Conceptualization, Project administration, Supervision, Writing—Review & Editing.
@@ Acknowledgments
Appendix A Full Results IBM Data
@@
In this part of the appendix, we provide the full results on the IBM datasets. These include all different patterns defined by Altman et al. [3]. The binary labels are defined using the threshold values 0.1, 0.2, 0.3, 0.5 and 0.9. The results on the HI-Small and LI-Large datasets are provided in Figure A.1 and Figure A.2, respectively.
@@
Figure A.1: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the HI-Small IBM dataset across all thresholds and patterns.
@@
Figure A.2: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the LI-Large IBM dataset across all thresholds and patterns.

### v2

@@ 5.1 Scalability
Figure 7 and Figure 7 hows the calculation times across the various network sizes of the synthetic data and IBM datasets, respectively. The time in Figure 7 is shown on a logarithmic scale since this provides a better distinction among the methods across dataset sizes. These results indicate that FlowScope is the fastest of all methods, although GARG-AML seems to also scale well to larger datasets, contrary to what we expected from the theoretical worst-case time complexity. In contrast, AutoAudit—although quite fast in the beginning—has difficulties with networks of the largest size. In addition, the distribution of the calculation time of AutoAudit is much wider and has more outliers, which indicates that it is highly sensitive to the network topology.
@@ 5.2.2 IBM Data
This section presents the results on the IBM datasets for the smurfing patterns, i.e., gather-scatter and scatter-gather, and the full ‘Is Laundering’ label. The thresholds applied for binary label definition are 0.1, 0.2 and 0.9. The results for the HI-Small and LI-Large datasets are presented in Figure 12 and Figure 13, respectively. The results for all patterns and thresholds are provided in Appendix B.
@@ 6.1 Limitations in the Data
GARG-AML assumes full visibility into all transactions, which is rarely the case in practice (Altman et al., 2023; Tariq & Hassani, 2025; Dumitrescu et al., 2022). Banks typically have access only to transactions that involve their own clients, including sender and receiver account numbers and client attributes. Transactions between two external accounts remain hidden, which makes second-order neighbourhoods incomplete. This can obscure smurfing structures or lead to incorrect classification (see Figure 14). Moreover, banks cannot link multiple third-party accounts that are owned by the same individual, thus further weakening pattern recognition.
Figure 14: Illustration of smurfing over different banks, which are represented by different colours. The (i)-icon indicates that the main (blue) bank has all client information about the corresponding account. The dotted lines indicate that the bank is unaware of these transactions/accounts. In most cases, the blue bank is aware of a big part of the transaction network.
This is mainly a limitation when only one bank would be guarding against money laundering. When two banks are involved, one of these two will have a clear picture of the transaction scheme. Although the transactions go through different banks in Figures 14(a) and (b), the blue bank still has the full transaction pattern. It becomes much harder to have a full picture when more than two banks are involved.
Given these limitations, we advocate for stronger collaboration both among banks and across countries. By enabling the secure sharing of transaction information, many of the current blind spots could be substantially reduced. In such an environment, GARG-AML would be positioned to deliver significantly greater value.
@@ 7 Conclusion
This paper introduces Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), which is a novel approach for detecting smurfing by quantifying the structural similarity of nodes’ neighbourhoods to known money laundering patterns via density-based features in the adjacency matrix. We validated GARG-AML on several synthetic datasets and compared it to state-of-the-art smurfing detection methods. The results of the experiments show that GARG-AML achieves strong detection power, with state-of-the-art performance of GARG-AML scores, and excellent scalability for large networks. Unlike approaches that are based on complex optimization techniques or machine learning models, GARG-AML relies on domain expertise and straightforward density-based calculations.
@@
Additionally, money laundering schemes typically exploit cross-bank and cross-border transaction chains, which fragment the available information and obscure detection efforts. This work advocates that enhanced collaboration, both at the national level among domestic banks and at the international level across jurisdictions, is essential. Establishing secure data-sharing frameworks and cooperative monitoring mechanisms would significantly strengthen the ability to trace complex laundering networks and improve the overall effectiveness of anti-money laundering measures.
Finally, detecting smurfs is an important research area for network security and intrusion detection (Yang et al., 2009; Shadrooh & Nørvåg, 2024). Further research could analyse whether similar featurisation of the patterns is possible for denial of service (DoS) or similar attacks.
@@
Appendix A Toy Example GARG-AML Scores
We demonstrate the calculations and results as described in Section 3, using a toy example. We start with a Watts-Strogatz network containing 20 nodes, each connecting to the next two nodes (initial degree equal to four), with a rewiring probability of $20\%$. We inject a smurfing pattern, where money originates at a new node and gets transferred to an existing node in the network using three mule accounts. The full network is visualised in Figure A.1.
We select two nodes to illustrate how the GARG-AML scores are calculated. We select node 23 because its second-order neighbourhood consitutes a pure smurfing pattern. We also select node 8, since its neighbours interact a lot with each other. So, the calculations for node 8 illustrate how the scores are impacted by these interactions.
Figure A.1: The undirected network with toy example. Nodes 20 to 23 are part of the infused smurfing pattern, and with node 19 the destination node.
A.1 Undirected Calculations
We start with the pure smurfing pattern using node 23. Its second-order neighbourhood is visualised in Figure A.2. Based on this neighbourhood, we construct the adjacency matrix as follows:
$\begin{array}[]{r}23\\ 19\\ 20\\ 21\\ 22\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&1&1&1\\ {\color[rgb]{0,1,0}1}&1&{\color[rgb]{1,0,0}0}&0&0\\ {\color[rgb]{0,1,0}1}&1&0&{\color[rgb]{1,0,0}0}&0\\ {\color[rgb]{0,1,0}1}&1&0&0&{\color[rgb]{1,0,0}0}\\ \end{pmatrix}$ (13)
The elements that should be zero, and hence are not taken into account, are indicated in red. Those elements that are certainly equal to one are indicated in green.
Figure A.2: Visualisation of the undirected second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
We calculate the scores by taking the upper-left block as block 1, the upper-right as block 2, and the lower-right as block 3. Using the notation as in Equations (5)-(8) , this results in the density calculations and sizes, where the size is the number of free elements:
$\text{score}_{1}$ $=$ $0$
$\text{score}_{2}$ $=$ $1$
$\text{score}_{3}$ $=$ $0$
$l_{1}$ $=$ $0$
$l_{2}$ $=$ $3$
$l_{3}$ $=$ $6$
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{score}_{2}-\frac{l_{1}\cdot\text{score}_{1}+l_{3}\cdot\text{score}_{3}}{l_{1}+l_{3}}$ (14)
$=$ $1$
As expected, the score that we assign to node 23 is 1. This indicates that its second-order neighbourhood forms a pure smurfing pattern.
The second-order neighbourhood of node 8 is shown in Figure A.3. We observe an interaction between node 7 and 9, which are part of the neighbours at distance one. Additionally, multiple interactions between nodes at distance two can be observed. This already indicates that the neighbourhood does not represent a pure smurfing pattern. The adjacency matrix is given in Equation (15), with the same colour scheme as before.
Figure A.3: Visualisation of the undirected second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
$\begin{array}[]{r}8\\ 3\\ 4\\ 5\\ 10\\ 11\\ 15\\ 6\\ 7\\ 9\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&1&1&0&0&1&0&0&1\\ {\color[rgb]{1,0,0}0}&1&{\color[rgb]{1,0,0}0}&1&1&0&0&0&1&0\\ {\color[rgb]{1,0,0}0}&1&1&{\color[rgb]{1,0,0}0}&0&0&0&0&1&0\\ {\color[rgb]{1,0,0}0}&0&1&0&{\color[rgb]{1,0,0}0}&1&0&1&0&0\\ {\color[rgb]{1,0,0}0}&0&0&0&1&{\color[rgb]{1,0,0}0}&0&1&0&0\\ {\color[rgb]{1,0,0}0}&1&0&0&0&0&{\color[rgb]{1,0,0}0}&0&0&1\\ {\color[rgb]{0,1,0}1}&0&0&0&1&1&0&{\color[rgb]{1,0,0}0}&0&1\\ {\color[rgb]{0,1,0}1}&0&1&1&0&0&0&0&{\color[rgb]{1,0,0}0}&0\\ {\color[rgb]{0,1,0}1}&1&0&0&0&0&1&1&0&{\color[rgb]{1,0,0}0}\end{pmatrix}$ (15)
The first row contains the node we will give a score, here node 8. The following rows represent the nodes at distance two, i.e., 3, 4, 5, 10, 11 and 15. The final rows represent the nodes at distance one, i.e., 6, 7 and 9.
This results in the following scores:
$\text{score}_{1}$ $=$ $0.4$
$\text{score}_{2}$ $=$ $0.33$
$\text{score}_{3}$ $=$ $0.33$
$l_{1}$ $=$ $30$
$l_{2}$ $=$ $18$
$l_{3}$ $=$ $6$
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{score}_{2}-\frac{l_{1}\cdot\text{score}_{1}+l_{3}\cdot\text{score}_{3}}{l_{1}+l_{3}}$ (16)
$=$ $-0.0556$
As indicated above, node 8 is not involved in a smurfing scheme. This is corroborated by the fact that the final score is negative.
A.2 Directed Calculations
We demonstrate the calculations of the directed GARG-AML scores on the same network as before, but with directed edges. This network is illustrated in Figure A.4. We again calculate the scores for node 23 and 8.
Figure A.4: The directed network with toy example. Nodes 20 to 23 are part of the infused smurfing pattern, and with node 19 the destination node.
The directed second-order neighbourhood for node 8 is shown in Figure A.5, and the corresponding adjacency matrix is provided in Equation (17). Note that we now have a different ordering of the nodes, using level 0 (node 23), level 1 (nodes 20, 21, 22) and level 2 (node 19), resulting in nine blocks, as defined in Section 3.3.
$\begin{array}[]{r}23\\ 20\\ 21\\ 22\\ 19\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&1&1&1&{\color[rgb]{1,0,0}0}\\ 0&{\color[rgb]{1,0,0}0}&0&0&1\\ 0&0&{\color[rgb]{1,0,0}0}&0&1\\ 0&0&0&{\color[rgb]{1,0,0}0}&1\\ {\color[rgb]{1,0,0}0}&0&0&0&{\color[rgb]{1,0,0}0}\end{pmatrix}$ (17)
Figure A.5: Visualisation of the directed second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
The density of the nine blocks is calculated. We will use the matrix representation as in Equation (11) to represent the densities of the blocks. For the adjacency matrix in Equation (17), this results in the following nine densities:
$\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 0&0&0\end{array}\right)$} (18)
The score for node 23 is then calculated as
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{mean}(\text{score}_{01},\text{score}_{12})$
$=$ $\text{mean}(1,1)-\text{mean}(0,0,0,0,0,0,0)$
$=$ $1.$ (20)
This again shows that the pure smurfing pattern gets a perfect score.
We do the same exercise for node 8. Here, level 0 consists of nodes 8 and 15, level 1 consists of nodes 6, 7 and 9, and level 2 consists of nodes 3, 4, 5, 10 and 11. Contrary to the second-order neighbourhood of node 23, we see on Figure A.6 that money flows in multiple directions, with nodes sending and receiving money to and from nodes across levels. This is already an indication that node 8 is not part of a smurfing scheme. The corresponding adjacency matrix is given in Equation (21).
Figure A.6: Visualisation of the directed second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
$\begin{array}[]{r}8\\ 15\\ 9\\ 6\\ 7\\ 3\\ 4\\ 5\\ 10\\ 11\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&1&0&0&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&0&0&0&0&0&0&0&0\\ 0&0&{\color[rgb]{1,0,0}0}&0&0&0&0&0&1&1\\ 1&0&0&{\color[rgb]{1,0,0}0}&0&0&0&0&0&0\\ 1&1&1&0&{\color[rgb]{1,0,0}0}&0&0&0&0&0\\ {\color[rgb]{1,0,0}0}&1&0&0&1&{\color[rgb]{1,0,0}0}&1&1&0&0\\ {\color[rgb]{1,0,0}0}&0&0&1&0&0&{\color[rgb]{1,0,0}0}&1&1&0\\ {\color[rgb]{1,0,0}0}&0&0&1&0&0&0&{\color[rgb]{1,0,0}0}&0&0\\ {\color[rgb]{1,0,0}0}&0&0&0&0&0&0&0&{\color[rgb]{1,0,0}0}&1\\ {\color[rgb]{1,0,0}0}&0&0&0&0&0&0&0&0&{\color[rgb]{1,0,0}0}\end{pmatrix}$ (21)
For this network, we get the following nine densities:
$\left(\begin{array}[]{ccc}0&0.1667&0\\ 0.5&0.1667&0.133\\ 0.2&0.2&0.25\end{array}\right)$} (22)
We observe that the densities are much different than for the pure smurfing pattern. This translates to the following score calculation:
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{mean}(\text{score}_{01},\text{score}_{12})$
$=$ $\text{mean}(0.1667,0.133)-\text{mean}(0,0,0.5,0.1667,0.2,0.2,0.25)$
$=$ $0.15-0.188$ (24)
$=$ $-0.038.$ (25)
The score for node 8 is below 0. Hence, we have no suspicion that this node is part of a smurfing scheme.
Appendix B Full Results IBM Data
@@
In this part of the appendix, we provide the full results on the IBM datasets. These include all different patterns defined by Altman et al. (2023). The binary labels are defined using the threshold values 0.1, 0.2, 0.3, 0.5 and 0.9. The results on the HI-Small and LI-Large datasets are provided in Figure B.1 and Figure B.2, respectively.
@@
Figure B.1: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the HI-Small IBM dataset across all thresholds and patterns.
@@
Figure B.2: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the LI-Large IBM dataset across all thresholds and patterns.